# Suicide Mortality in Spain (2010-2022): Temporal Trends, Spatial Patterns, and Risk Factors

### v1

@@ 1 Introduction
According to the most recent report from the World Health Organization, more than 700,000 people died by suicide worldwide in 2021 (World Health Organization, , 2025), accounting for approximately 1.1% of all global deaths. The global age-standardized suicide rate was notably higher in males than in females (12.3 and 5.6 per 100,000 inhabitants, respectively) with the male-to-female suicide ratio reaching 3.2 in high-income countries. Although these countries show the highest age-standardized suicide rates, it is concerning that nearly 75% of the total number of suicides occur in low- and middle-income countries. These figures provide an overall picture, yet suicide rate trends differ markedly between regions. While suicide mortality has decreased in several parts of Europe, the Americas, and Australasia since the 2000s, countries such as the UK, Brazil, Mexico, the USA, South Korea, and Australia have experienced rising trends (Alicandro et al., , 2019). Among these, the United States stands out as a particularly concerning case. Suicide remains a leading cause of death in the country and has shown a troubling upward trend in recent years. Between 2020 and 2023, the national suicide rate increased by 4.5%. In 2023 alone, more than 49,000 lives were lost to suicide, corresponding to an age-standardized rate of 14.1 per 100,000 inhabitants. Additionally, an estimated 12.8 million adults seriously considered ending their lives, 3.7 million made a plan, and 1.5 million attempted suicide (Centers for Disease Control and Prevention, , 2025). In Europe, from 2000 to 2019, age-standardized suicide mortality rates declined by 2.3% in females and 3.4% in males (Ilic and Ilic, , 2022). However, recent studies indicate that some countries (including Iceland, Norway, Sweden, and Spain among others) have experienced an upward trend in suicide rates over the past decade (Zuin and de Leo, , 2025). Beyond the pronounced sex differences, age is a well-recognized risk factor for suicide, with older adults generally exhibiting higher mortality rates across most countries (World Health Organization, , 2014). This increased risk is especially notable among individuals suffering from physical illnesses, depression, and anxiety, conditions that frequently contribute to suicidal behavior in this population (Conwell et al., , 2011). Consistent with these patterns, estimates from the most recent Global Burden of Disease Study (Weaver et al., , 2025) indicate that the mean age at the time of death from suicide has increased over the past three decades, rising by four years among males (from 43.0 to 47.0) and by five years among females (from 41.9 to 46.9).
@@
To better understand the dynamics and determinants of suicide mortality, numerous studies have examined its temporal evolution (often stratified by sex and age groups) or have sought to identify spatial patterns that highlight inequalities across subnational administrative units. Depending on their objectives and data availability, these studies typically adopt temporal, spatial, or spatio-temporal modelling frameworks. When analysing temporal changes in national suicide mortality trends, researchers have employed a variety of methods, ranging from simpler statistical techniques such as moving average smoothing (Ivey-Stephenson, , 2017) and regression models for change-point analysis (Pompili et al., , 2009; Hofstra et al., , 2018; Ramalle-Gómara et al., , 2023), to more advanced approaches like age-period-cohort (APC) models (Granizo et al., , 1996; Riebler et al., , 2012), which are particularly useful for identifying long-term patterns. Beyond purely temporal approaches, an increasing number of studies have investigated the spatial and spatio-temporal distribution of suicide mortality rates, aiming to identify regional disparities and examine the influence of socioeconomic factors on geographic disparities. Methodological approaches in this area range from descriptive mapping (Qi et al., , 2014) to the application of inferential spatial models, including Poisson regressions with spatially structured random effects (Middleton et al., , 2006; Salmeron et al., , 2013) and the use of space-time scan statistics to detect potential high-risk clusters (Núñez-González et al., , 2018).
@@
In this paper, we analyze annual suicide mortality data from Spanish provinces, disaggregated by sex and age group. This detailed stratification allows us to explore temporal trends and spatial distribution patterns across different population groups. To this end, we apply Poisson mixed models separately for each sex, incorporating structured random effects that borrow information from neighboring units to obtain reliable mortality rate estimates by effectively capturing spatial, temporal, and age-group dependencies (Goicoa et al., , 2016, 2019). Additionally, we include the percentage of the population living in rural areas, the unemployment rate, and the at-risk-of-poverty rate as explanatory variables to assess their influence on suicide mortality rates and to gain deeper insight into the factors driving geographic and temporal disparities. These variables are modelled using an ecological regression framework to assess area-level associations while accounting for potential spatial confounding (Urdangarin et al., , 2024). In summary, our research focuses on two main objectives: (i) analyzing sex-specific temporal and spatial patterns in age-stratified suicide mortality rates in Spanish provinces, and (ii) examining the influence of socioeconomic and contextual factors on suicide rates while accounting for potential confounding issues.
@@ 2 Data
This study uses suicide mortality counts and corresponding population data from the Spanish Statistical Office (INE) for the period 2010-2022. The data, which includes sex- and age-specific counts (ICD-10 codes X60-X84), is stratified by year and province across continental Spain. To assess potential associations with suicide mortality, we also include several socioeconomic and contextual variables. These consist of two spatial variables, the percentage of the population living in rural areas and the provincial unemployment rate, and two temporal variables, the national annual employment rate and the at-risk-of-poverty rate. To construct a provincial-level rurality indicator, we use data from the Geographic Information System of the European Commission (GISCO-Eurostat, , 2024), which classifies local administrative units (municipalities in the case of Spain) into three categories of urbanisation: cities, towns or suburbs, and rural areas. Combining the 2020 classification with Spanish population data, we calculate the percentage of the population living in rural areas for each province. We obtain the remaining covariates (i.e., the provincial unemployment rate, national annual employment rate, and at-risk-of-poverty rate) from the Spanish Statistical Office.
@@
In 2022, suicide remained the leading cause of external mortality in Spain, accounting for 22.8% of all such deaths, ahead of unintentional threats to breathing (22.1%), accidental falls (20.4%), and transport accidents (9.8%) (Instituto Nacional de Estadística, , 2023). Between 2010 and 2022, an average of 3,639 suicides were reported annually in Spain, representing 0.9% of all deaths. The burden of suicide is highest among younger age groups. For individuals under 25, suicide accounts for 5.9% of all deaths. This proportion rises to 11.7% in the 25-44 age group, then decreases to 2.8% for those aged 45-64 and 0.3% for individuals aged 65 and over. Our analysis focuses on the provinces of continental Spain, excluding the Balearic and Canary Islands and the autonomous cities of Ceuta and Melilla. A map of the provinces is provided in Supplementary Figure S1. A total of 43,604 suicide deaths were recorded during this period, with 74.9% correspond to males and 25.1% to females. This distribution aligns with the pattern observed across the European Union, where males account for 76.7% of all suicides (Eurostat, , 2024).
@@
In an initial exploratory analysis, we identified potential under-reporting in Madrid’s data for the years 2010, 2011, and 2012. Accordingly, these values were treated as missing and we then compute the national suicide mortality rates for those years by excluding both the population and mortality data from Madrid. Furthermore, when aggregating data into ten-year age intervals, suicide deaths in the youngest group (0-10 years) were negligible, and this group was therefore excluded from the analysis. Figure 1 illustrates the average annual number of suicide deaths and crude mortality rates by ten-year age groups and sex in continental Spain from 2010 to 2022. The figure shows that males consistently exhibit higher mortality rates than females across all age groups. Over the study period, the average suicide mortality rate among males is 13.4 per 100,000 inhabitants, about three times higher than for females (4.3 per 100,000). Age-specific trends reveal sex differences. In males, suicide mortality rates increase, peaking in the 50-59 age group before a slight decline in the 60-69 age group, followed by a steady rise with advancing age. Conversely, females show a similar peak around 50-59, but their rates remain relatively stable in subsequent age groups. These patterns underscore the strong influence of both sex and age on suicide mortality rates.
@@
Figure 1: Average annual number of suicide deaths and crude mortality rates by age group and sex in continental Spain, 2010-2022.
@@
Figure 2: Maps of suicide mortality rates by province in continental Spain, 2010-2022.
@@
Figure 3: Annual suicide mortality rates by sex and age groups. The dashed black line indicates the overall mortality rate.
Table 1 presents suicide mortality rates (with 95% confidence intervals in brackets) disaggregated by sex, age groups, and rurality level. The rurality variable classifies the percentage of the population residing in rural areas within each province into three categories: low (less than 20%), medium (20-40%), and high (over 40%). We compute confidence intervals using a normal approximation under the assumption that suicide counts follow a Poisson distribution. Across all age groups, suicide mortality rates are consistently higher in provinces with medium or high rurality levels, with the exception of female mortality in the 40-70 age group. These differences are particularly pronounced among males, with rates in highly rural areas exceeding those in low-rurality provinces by more than 30%. In contrast, female rates remain relatively stable across rurality levels, except for a slight increase among those aged 70 and above.
Table 1: Suicide mortality rates (per 100,000 inhabitants), with 95% confidence intervals in brackets, by sex, age groups, and rurality level in Spanish provinces, 2010-2022.
@@
3 Methods
Given the pronounced differences observed in suicide mortality rates between males and females in Spain, we fitted separate models for each sex. This was necessary to accurately capture distinct age, spatial, and temporal trends. As the previous descriptive analysis showed, the observed patterns demonstrate marked sex-specific dynamics, meaning a shared spatial or temporal correlation structure could not be assumed across both sexes.
Initially, we aim to fit sex-specific models that incorporate structured random effects (including their interactions) for age group, province, and year, to estimate the spatio-temporal evolution of suicide mortality rates while sharing information across adjacent age groups. We found that fully stratified age-space-time models exhibited excessive smoothing. This was likely due to the limited number of suicide cases within each stratum, a problem particularly pronounced among females. Such excessive smoothing, which hinders the models’ ability to capture meaningful variation, is a documented limitation in the literature. For example, Retegui et al., (2025) have shown that spatial analyses in very small areas can lead to this effect. Fitting separate spatio-temporal models for each age group is also disadvantageous, as it fails to overcome data sparsity and prevents the model from leveraging shared patterns across age groups. To address these limitations and avoid over-smoothing, we decided to model the age-time and age-space dimensions separately, enabling a more reliable estimation of patterns within each component.
3.1 Age-time interaction models
First, we analyze the data using a Bayesian hierarchical model that incorporates an age-time interaction effect. Let $O_{at}$ denote the observed number of suicide deaths in age group $a=1,\dots,8$ (corresponding to ages 10-19, …, 70-79, and 80+), and time period $t=1,\dots,13$, representing years 2010 to 2022. Let $N_{at}$ and $\lambda_{at}$ denote the population at risk and the suicide mortality rate, respectively, for age group $a$ in year $t$. Conditional on the rates, we assume that the number of observed cases follows a Poisson distribution with mean $\mu_{at}=N_{at}\lambda_{at}$, that is
Then, the log-rate is modelled as
$\log{\lambda_{at}}=\alpha+\phi_{a}+\gamma_{t}+\delta_{at},$ (2)
The model performance led us to select a first-order random walk (RW1) prior for the age group random effect, ${\mbox{\boldmath{$\phi$}}}=(\phi_{1},\ldots,\phi_{8})^{{}^{\prime}}$. We chose a RW1 prior for the temporal effect ${\mbox{\boldmath{$\gamma$}}}=(\gamma_{1},\ldots,\gamma_{13})^{{}^{\prime}}$ in the male model, but we applied a second-order random walk (RW2) prior for females. The RW2 prior introduces a higher degree of smoothness by borrowing strength from second-order temporal neighbors, which is particularly beneficial given the lower number of suicide cases observed among females. So, we assume that
$\mbox{\boldmath{$\phi$}}\sim N({\bf 0},[\tau_{\phi}{\bf R}_{\phi}]^{-}),\qquad\mbox{and}\qquad\mbox{\boldmath{$\gamma$}}\sim N({\bf 0},[\tau_{\gamma}{\bf R}_{\gamma}]^{-}),$
where the symbol {-} denotes the Moore-Penrose generalized inverse of a matrix, $\tau_{\phi}$ and $\tau_{\gamma}$ are precision parameters, and ${\bf R}_{\phi}$ and ${\bf R}_{\gamma}$ are structure matrices that define the precision structure of the RW priors, inducing smoothness across adjacent age groups and time periods, respectively. Finally, we assume a multivariate normal prior for the age-time random effect, $\mbox{\boldmath{$\delta$}}\sim N({\bf 0},[\tau_{\delta}{\bf R}_{\delta}]^{-})$ where ${\bf R}_{\delta}$ is the precision matrix obtained as the Kronecker product corresponding to the four different types of interactions originally proposed by Knorr-Held, (2000). A detailed summary of the prior assumptions for each random effect is provided in Table 2. See Goicoa et al., (2018) for a discussion on identifiability constraints in these models and their implementation using INLA.
Age-time models Age-space models
Age effect (RW1) $\mbox{\boldmath{$\phi$}}\sim N({\bf 0},[\tau_{\phi}{\bf R}_{\phi}]^{-})$
Temporal effect (RW1/RW2) $\mbox{\boldmath{$\gamma$}}\sim N({\bf 0},[\tau_{\gamma}{\bf R}_{\gamma}]^{-})$ -
Spatial effect (iCAR) - $\mbox{\boldmath{$\xi$}}\sim N({\bf 0},[\tau_{\xi}{\bf R}_{\xi}]^{-})$
Interaction effect $\mbox{\boldmath{$\delta$}}\sim N({\bf 0},[\tau_{\delta}{\bf R}_{\delta}]^{-})$ $\mbox{\boldmath{$\delta$}}\sim N({\bf 0},[\tau_{\delta}{\bf R}_{\delta}]^{-})$
3.2 Age-space interaction models
Next, we analyze the data by fitting a Bayesian hierarchical model that incorporates an age-space interaction effect, allowing us to capture spatial variation in suicide mortality rates across provinces while accounting for age-specific differences. Let $O_{as}$ denote the observed number of suicide deaths in age group $a=1,\dots,8$ (corresponding to ages 10-19, …, 70-79, and 80+), and small area $s=1,\dots,47$, representing the provinces of continental Spain. Let $N_{as}$ and $\lambda_{as}$ denote the population at risk and the suicide mortality rate, respectively, for age group $a$ in province $s$. As in the previous models, we assume that
Then, the log-rate is modelled as
$\log{\lambda_{as}}=\alpha+\phi_{a}+\xi_{s}+\delta_{as},$ (4)
where $\alpha$ denotes the overall intercept (baseline log-rate), $\phi_{a}$ and $\xi_{s}$ represents the main effects for age group and space, respectively, and $\delta_{as}$ denotes the age-space interaction term. For these particular set of models, we selected a RW1 prior distribution for age-group random effect ${\mbox{\boldmath{$\phi$}}}=(\phi_{1},\ldots,\phi_{8})^{{}^{\prime}}$ and an intrinsic conditional autoregressive (iCAR) prior for the spatial random effect ${\mbox{\boldmath{$\xi$}}}=(\xi_{1},\ldots,\xi_{47})^{{}^{\prime}}$. That is,
$\mbox{\boldmath{$\phi$}}\sim N({\bf 0},[\tau_{\phi}{\bf R}_{\phi}]^{-}),\qquad\mbox{and}\qquad\mbox{\boldmath{$\xi$}}\sim N({\bf 0},[\tau_{\xi}{\bf R}_{\xi}]^{-}),$
The restricted spatial regression (RSR) method is probably one of the most widely adopted strategies to handle spatial confounding (Reich et al., , 2006; Hodges and Reich, , 2010), although several alternative approaches have been proposed. Recent studies, however, argue that RSR implicitly assumes the absence of unobserved covariates correlated with the observed ones, since it constrains the random effects to be orthogonal to the fixed effects (Gilbert et al., , 2024). In contrast, the spatial+ approach proposed by Dupont et al., (2022) mitigates spatial confounding by removing the spatial dependence of the covariates. In this study, we implement the simplified spatial+ method introduced by Urdangarin et al., (2024), a modification that avoids separately fitting spatial models for the covariates to eliminate their spatial dependence. Below, we provide a brief overview of this method.
${\bf X}=a_{1}{\bf U}_{1}+\ldots+a_{n}{\bf U}_{n}.$
3.4 Model implementation
The models described in the previous sections are implemented within a fully Bayesian framework using the integrated nested Laplace approximation (INLA) method (Rue et al., , 2009), a widely used technique for approximate Bayesian inference in applied statistics (Rue et al., , 2017). Specifically, we fit the models using the novel hybrid approximation approach that combines the Laplace method with a low-rank Variational Bayes correction to the posterior mean (van Niekerk et al., , 2023; van Niekerk and Rue, , 2024) using the R-INLA stable version 25.06.07 on R-4.5.1.
Regarding model hyperparameters, we assign improper uniform priors to all standard deviations (i.e., the square roots of the inverse precision parameters), while we use weakly informative normal priors with mean zero and precision 0.001 for the fixed effects. To ensure meaningful comparison of precision parameters within each model, we scaled all structure matrices such that the geometric mean of their marginal variances equals one (Sørbye and Rue, , 2014). The R code to reproduce the all the results is available at https://github.com/spatialstatisticsupna/Suicides_Spain.
4 Results
We first present key results from the interaction models used to analyze age-group, temporal, and spatial patterns in suicide mortality rates among males and females. We then assess the influence of explanatory variables on the temporal trends and geographical distribution of these rates.

### v2

@@ Abstract
Background: Suicide remains a major public health concern worldwide, responsible for more than 700,000 deaths in 2021, accounting for approximately 1.1% of all global deaths. While many high-income countries have reported declines in age-standardized suicide rates over the past two decades, recent evidence from Spain indicates increasing mortality among women, whereas suicide rates among men have remained relatively stable. To better understand these patterns and their potential underlying determinants, this study examines the spatial and temporal patterns of age-stratified suicide mortality across Spanish provinces from 2010 to 2022, with particular attention to sex-specific differences.
Methods: Mixed Poisson models were applied to analyze provincial- and temporal-level suicide mortality rates, stratified by age and sex. The models accounted for spatial and temporal confounding effects and examined associations with various socioeconomic and contextual factors, including rurality and unemployment.
@@
Conclusions: The use of sophisticated statistical models permits the detection of underlying patterns, revealing both geographic and temporal disparities in suicide mortality across Spanish provinces.
@@ 1 Introduction
According to the most recent report from the World Health Organization, more than 700,000 people died by suicide worldwide in 2021 (world2025suicide), accounting for approximately 1.1% of all global deaths. The global age-standardized suicide rate was notably higher in males than in females (12.3 and 5.6 per 100,000 inhabitants, respectively) with the male-to-female suicide ratio reaching 3.2 in high-income countries. Although these countries show the highest age-standardized suicide rates, it is concerning that nearly 75% of the total number of suicides occur in low- and middle-income countries. These figures provide an overall picture, yet suicide rate trends differ markedly between regions. While suicide mortality has decreased in several parts of Europe, the Americas, and Australasia since the 2000s, countries such as the UK, Brazil, Mexico, the USA, South Korea, and Australia have experienced rising trends (alicandro2019worldwide). Among these, the United States represents a notable example, with suicide remaining one of the leading causes of death and exhibiting a gradual increase in recent years. Between 2020 and 2023, the national suicide rate increased by 4.5%. In 2023 alone, more than 49,000 lives were lost to suicide, corresponding to an age-standardized rate of 14.1 per 100,000 inhabitants. Additionally, an estimated 12.8 million adults seriously considered ending their lives, 3.7 million made a plan, and 1.5 million attempted suicide (CDC2025). In Europe, from 2000 to 2019, age-standardized suicide mortality rates declined by 2.3% in females and 3.4% in males (ilic2022worldwide). However, recent studies have reported some variation in suicide rates over the past decade in several countries, including Iceland, Norway, Sweden, and Spain (zuin2025suicide), although in many cases these changes are not statistically significant.
Beyond the pronounced sex differences, age is a well-recognized factor in suicide risk, with older adults generally exhibiting higher mortality rates across most countries (WHO2014). However, the relationship between age and suicide is complex and often non-linear, with specific age groups exhibiting higher vulnerability depending on the social context. This increased risk is especially notable among individuals suffering from physical illnesses, affective disorders and other mental disorders, conditions that frequently contribute to suicidal behavior (conwell2011suicide; kim2025risk). Consistent with these patterns, estimates from the most recent Global Burden of Disease Study (weaver2025global) indicate that the mean age at the time of death from suicide has increased over the past three decades, rising by four years among males (from 43.0 to 47.0) and by five years among females (from 41.9 to 46.9).
@@
To better understand the dynamics and determinants of suicide mortality, numerous studies have examined its temporal evolution (often stratified by sex and age groups) or have sought to identify spatial patterns that highlight inequalities across subnational administrative units. Depending on their objectives and data availability, these studies typically adopt temporal, spatial, or spatio-temporal modelling frameworks. When analysing temporal changes in national suicide mortality trends, researchers have employed a variety of methods, ranging from simpler statistical techniques such as moving average smoothing (ivey2017suicide) and regression models for change-point analysis (pompili2009suicide; hofstra2018springtime; ramalle2023trends), to more advanced approaches like age-period-cohort models (granizo1996age; riebler2012gender), which are particularly useful for identifying long-term patterns. Beyond purely temporal approaches, an increasing number of studies have investigated the spatial and spatio-temporal distribution of suicide mortality rates, aiming to identify regional disparities and examine the influence of socioeconomic factors on geographic disparities. Methodological approaches in this area range from descriptive mapping (qi2014dynamic) to the application of inferential spatial models, including Poisson regressions with spatially structured random effects (middleton2006geography; salmeron2013time) and the use of space-time scan statistics to detect potential high-risk clusters (nunez2018trends).
In this paper, we analyze annual suicide mortality in Spain from 2010 to 2022, using data disaggregated by sex and age group at the provincial level. This detailed stratification allows us to examine temporal trends and spatial distribution patterns across different population groups. Previous studies have reported a marked increase in suicide mortality among Spanish women between 2010 and 2016, while rates among men have remained relatively stable (cayuela2020suicide). The study period also includes the COVID-19 pandemic, which exacerbated several known risk factors for suicide and may have influenced changes in suicide mortality by gender, age group, and geographic region (lantos2024impact). Accordingly, this study provides an opportunity to evaluate the potential impact of this crisis on suicide mortality in Spain. Furthermore, several studies have highlighted pronounced geographical heterogeneity in suicide mortality across the country, with higher rates typically observed in regions with lower socioeconomic development (santurtun2018does).
@@
To this end, we apply Poisson mixed models separately for each sex, incorporating structured random effects that borrow information from neighboring units to obtain reliable mortality rate estimates by effectively capturing spatial, temporal, and age-group dependencies (goicoa2016age; goicoa2019flexible). Additionally, we include the percentage of the population living in rural areas, the unemployment rate, and the at-risk-of-poverty rate as explanatory variables to assess their influence on suicide mortality rates and to gain deeper insight into the factors driving geographic and temporal disparities. These variables are modelled using an ecological regression framework to assess area-level associations while accounting for potential spatial confounding (urdangarin2024simplified). In summary, our research focuses on two main objectives: (i) analyzing sex-specific temporal and spatial patterns in age-stratified suicide mortality rates in Spanish provinces, and (ii) examining the influence of socioeconomic and contextual factors on suicide rates while accounting for potential confounding issues.
@@ 2 Data
This study uses suicide mortality counts and corresponding population data from the Spanish National Statistics Institute (INE) for the period 2010–2022. The data, which includes sex- and age-specific counts (ICD-10 codes X60–X84), is stratified by year and province across continental Spain. To assess potential associations with suicide mortality, we also include several socioeconomic and contextual variables. These consist of two spatial variables, the percentage of the population living in rural areas and the provincial unemployment rate, and two temporal variables, the national annual employment rate and the at-risk-of-poverty rate. To construct a provincial-level rurality indicator, we use data from the Geographic Information System of the European Commission (eurostat2024gisco), which classifies local administrative units (municipalities in the case of Spain) into three categories of urbanisation: cities, towns or suburbs, and rural areas. Combining the 2020 classification with Spanish population data, we calculate the percentage of the population living in rural areas for each province. We obtain the remaining covariates (i.e., the provincial unemployment rate, national annual employment rate, and at-risk-of-poverty rate) from the National Statistics Institute.
@@
In 2022, suicide remained the leading cause of external mortality in Spain, accounting for 22.8% of all such deaths, ahead of unintentional threats to breathing (22.1%), accidental falls (20.4%), and transport accidents (9.8%) (INE2023defunciones). Between 2010 and 2022, an average of 3,639 suicides were reported annually in Spain, representing 0.9% of all deaths. The burden of suicide is highest among younger age groups. For individuals under 25, suicide accounts for 5.9% of all deaths. This proportion rises to 11.7% in the 25–44 age group, then decreases to 2.8% for those aged 45–64 and 0.3% for individuals aged 65 and over. Our analysis focuses on the provinces of continental Spain, excluding the Balearic and Canary Islands and the autonomous cities of Ceuta and Melilla. A map of the provinces is provided in Supplementary Figure S1. A total of 43,604 suicide deaths were recorded during this period, with 74.9% correspond to males and 25.1% to females. This distribution aligns with the pattern observed across the European Union, where males account for 76.7% of all suicides (Eurostat2024).
3 Methods
Given the pronounced differences in suicide mortality rates between males and females in Spain, we fitted separate models for each sex. This approach allows for a more accurate representation of distinct age, spatial, and temporal trends. Because of these marked sex-specific dynamics, a shared spatial or temporal correlation structure could not be assumed across both sexes.
Initially, we aim to fit sex-specific models that incorporate structured random effects (including their interactions) for age group, province, and year, to estimate the spatio-temporal evolution of suicide mortality rates while sharing information across adjacent age groups. We found that fully stratified age-space-time models exhibited excessive smoothing. This was likely due to the limited number of suicide cases within each stratum, a problem particularly pronounced among females. Such excessive smoothing, which hinders the models’ ability to capture meaningful variation, is a documented limitation in the literature. For example, retegui2025prior have shown that spatial analyses in very small areas can lead to this effect. Fitting separate spatio-temporal models for each age group is also disadvantageous, as it fails to overcome data sparsity and prevents the model from leveraging shared patterns across age groups. To address these limitations and avoid over-smoothing, we decided to model the age-time and age-space dimensions separately, enabling a more reliable estimation of patterns within each component.
Given the structure of the age-space interaction models applied in Section 3.2, we restricted the analysis to the provinces of continental Spain. Including the islands would have required artificially connecting them to the mainland in the spatial adjacency graph, which could have biased the spatial estimates by not reflecting true socioeconomic or health relationships.
3.1 Age-time interaction models
First, we analyze the data using a Bayesian hierarchical model that incorporates an age-time interaction effect. Let $O_{at}$ denote the observed number of suicide deaths in age group $a=1,\dots,8$ (corresponding to ages 10–19, …, 70–79, and 80+), and time period $t=1,\dots,13$, representing years 2010 to 2022. Let $N_{at}$ and $\lambda_{at}$ denote the population at risk and the suicide mortality rate, respectively, for age group $a$ in year $t$. Conditional on the rates, we assume that the number of observed cases follows a Poisson distribution with mean $\mu_{at}=N_{at}\lambda_{at}$, that is
Then, the log-rate is modelled as
$\log{\lambda_{at}}=\alpha+\phi_{a}+\gamma_{t}+\delta_{at},$ (2)
The model performance led us to select a first-order random walk (RW1) prior for the age group random effect, ${\mbox{{$\phi$}}}=(\phi_{1},\ldots,\phi_{8})^{{}^{\prime}}$. We chose a RW1 prior for the temporal effect ${\mbox{{$\gamma$}}}=(\gamma_{1},\ldots,\gamma_{13})^{{}^{\prime}}$ in the male model, but we applied a second-order random walk (RW2) prior for females. The RW2 prior introduces a higher degree of smoothness by borrowing strength from second-order temporal neighbors, which is particularly beneficial given the lower number of suicide cases observed among females. So, we assume that
$\mbox{{$\phi$}}\sim N({\bf 0},[\tau_{\phi}{\bf R}_{\phi}]^{-}),\qquad\mbox{and}\qquad\mbox{{$\gamma$}}\sim N({\bf 0},[\tau_{\gamma}{\bf R}_{\gamma}]^{-}),$
where the symbol {-} denotes the Moore-Penrose generalized inverse of a matrix, $\tau_{\phi}$ and $\tau_{\gamma}$ are precision parameters, and ${\bf R}_{\phi}$ and ${\bf R}_{\gamma}$ are structure matrices that define the precision structure of the RW priors, inducing smoothness across adjacent age groups and time periods, respectively. Finally, we assume a multivariate normal prior for the age-time random effect, $\mbox{{$\delta$}}\sim N({\bf 0},[\tau_{\delta}{\bf R}_{\delta}]^{-})$ where ${\bf R}_{\delta}$ is the precision matrix obtained as the Kronecker product (denoted by the symbol $\otimes$) corresponding to the four different types of interactions originally proposed by knorr2000. A detailed summary of the prior assumptions for each random effect is provided in Table 1. See goicoa2018spatio for a discussion on identifiability constraints in these models and their implementation using INLA.
3.2 Age-space interaction models
Next, we analyze the data by fitting a Bayesian hierarchical model that incorporates an age-space interaction effect, allowing us to capture spatial variation in suicide mortality rates across provinces while accounting for age-specific differences. Let $O_{as}$ denote the observed number of suicide deaths in age group $a=1,\dots,8$ (corresponding to ages 10–19, …, 70–79, and 80+), and small area $s=1,\dots,47$, representing the provinces of continental Spain. Let $N_{as}$ and $\lambda_{as}$ denote the population at risk and the suicide mortality rate, respectively, for age group $a$ in province $s$. As in the previous models, we assume that
Then, the log-rate is modelled as
$\log{\lambda_{as}}=\alpha+\phi_{a}+\xi_{s}+\delta_{as},$ (4)
where $\alpha$ denotes the overall intercept (baseline log-rate), $\phi_{a}$ and $\xi_{s}$ represents the main effects for age group and space, respectively, and $\delta_{as}$ denotes the age-space interaction term. For these particular set of models, we selected a RW1 prior distribution for age-group random effect ${\mbox{{$\phi$}}}=(\phi_{1},\ldots,\phi_{8})^{{}^{\prime}}$ and an intrinsic conditional autoregressive (iCAR) prior for the spatial random effect ${\mbox{{$\xi$}}}=(\xi_{1},\ldots,\xi_{47})^{{}^{\prime}}$. That is,
$\mbox{{$\phi$}}\sim N({\bf 0},[\tau_{\phi}{\bf R}_{\phi}]^{-}),\qquad\mbox{and}\qquad\mbox{{$\xi$}}\sim N({\bf 0},[\tau_{\xi}{\bf R}_{\xi}]^{-}),$
Age-time models Age-space models
Age effect (RW1) $\mbox{{$\phi$}}\sim N({\bf 0},[\tau_{\phi}{\bf R}_{\phi}]^{-})$
Temporal effect (RW1/RW2) $\mbox{{$\gamma$}}\sim N({\bf 0},[\tau_{\gamma}{\bf R}_{\gamma}]^{-})$ -
Spatial effect (iCAR) - $\mbox{{$\xi$}}\sim N({\bf 0},[\tau_{\xi}{\bf R}_{\xi}]^{-})$
Interaction effect $\mbox{{$\delta$}}\sim N({\bf 0},[\tau_{\delta}{\bf R}_{\delta}]^{-})$ $\mbox{{$\delta$}}\sim N({\bf 0},[\tau_{\delta}{\bf R}_{\delta}]^{-})$
The restricted spatial regression (RSR) method is probably one of the most widely adopted strategies to handle spatial confounding (reich2006effects; hodges2010adding), although several alternative approaches have been proposed. Recent studies, however, argue that RSR implicitly assumes the absence of unobserved covariates correlated with the observed ones, since it constrains the random effects to be orthogonal to the fixed effects (gilbert2024). In contrast, the spatial+ approach proposed by dupont2022spatial+ mitigates spatial confounding by removing the spatial dependence of the covariates. In this study, we implement the simplified spatial+ method introduced by urdangarin2024simplified, a modification that avoids separately fitting spatial models for the covariates to eliminate their spatial dependence. Below, we provide a brief overview of this method.
${\bf X}=a_{1}{\bf U}_{1}+\ldots+a_{n}{\bf U}_{n}.$
3.4 Model implementation
The models described in the previous sections are implemented within a fully Bayesian framework using the integrated nested Laplace approximation (INLA) method (rue2009approximate), a widely used technique for approximate Bayesian inference in applied statistics (rue2017bayesian). Specifically, we fit the models using the novel hybrid approximation approach that combines the Laplace method with a low-rank Variational Bayes correction to the posterior mean (van2023new; van2024low) using the R-INLA stable version 25.06.07 on R-4.5.1.
Regarding model hyperparameters, we assign improper uniform priors to all standard deviations (i.e., the square roots of the inverse precision parameters), while we use weakly informative normal priors with mean zero and precision 0.001 for the fixed effects. To ensure meaningful comparison of precision parameters within each model, we scaled all structure matrices such that the geometric mean of their marginal variances equals one (sorbye2014scaling).
4 Results
We first describe the main descriptive results of suicide mortality across sex, age groups, and provinces to provide context for the modelling analyses. We then report key results from the interaction models examining age, temporal, and spatial patterns in suicide mortality among males and females. Finally, we assess the influence of explanatory variables on the temporal trends and geographical distribution of these rates.
4.1 Descriptive analysis
@@
In an initial exploratory analysis, we identified potential under-reporting in Madrid’s data for the years 2010, 2011, and 2012. This finding aligns with explanations by some authors (SANTURTUN2022150), who note that until 2013, the INE (National Statistics Institute) lacked access to data from Madrid’s Forensic Anatomical Institute. Accordingly, these values were treated as missing and we then compute the national suicide mortality rates for those years by excluding both the population and mortality data from Madrid. Furthermore, when aggregating data into ten-year age intervals, suicide deaths in the youngest group (0–10 years) were negligible, and this group was therefore excluded from the analysis. Figure 1 illustrates the average annual number of suicide deaths and crude mortality rates by ten-year age groups and sex in continental Spain from 2010 to 2022. The figure shows that males consistently exhibit higher mortality rates than females across all age groups. Over the study period, the average suicide mortality rate among males is 13.4 per 100,000 inhabitants, about three times higher than for females (4.3 per 100,000). Age-specific trends reveal sex differences. In males, suicide mortality rates increase, peaking in the 50–59 age group before a slight decline in the 60–69 age group, followed by a steady rise with advancing age. Conversely, females show a similar peak around 50–59, but their rates remain relatively stable in subsequent age groups. These patterns underscore the strong influence of both sex and age on suicide mortality rates.
@@
Figure 1: Average annual number of suicide deaths and crude mortality rates by age group and sex in continental Spain, 2010–2022.
@@
Figure 2: Maps of suicide mortality rates (per 100,000 inhabitants) by province in continental Spain, 2010–2022. (a) Males; (b) Females.
@@
Figure 3: Annual suicide mortality rates by sex and age groups. The dashed black line indicates the overall mortality rate. (a) Male population; (b) Female population.
Table 2 presents suicide mortality rates (with 95% confidence intervals in brackets) disaggregated by sex, age groups, and rurality level. The rurality variable classifies the percentage of the population residing in rural areas within each province into three categories: low (less than 20%), medium (20–40%), and high (over 40%). We compute confidence intervals using a normal approximation under the assumption that suicide counts follow a Poisson distribution. Across all age groups, suicide mortality rates are consistently higher in provinces with medium or high rurality levels, with the exception of female mortality in the 40–70 age group. These differences are particularly pronounced among males, with rates in highly rural areas exceeding those in low-rurality provinces by more than 30%. In contrast, female rates remain relatively stable across rurality levels, except for a slight increase among those aged 70 and above.
Table 2: Suicide mortality rates (per 100,000 inhabitants), with 95% confidence intervals in brackets, by sex, age groups, and rurality level in Spanish provinces, 2010–2022.